The manager at a store selling tea plans to mix a more expensive tea that costs $5 dollars per pound with a less expensive tea that costs $3 a pound to create a 100 pound blend that will sell for $4.50 per pound. How many pounds of each type of tea are required?
Answer by mananth(16946)

The manager at a store selling tea plans to mix a more expensive tea that costs $5 dollars per pound with a less expensive tea that costs $3 a pound to create a 100 pound blend that will sell for $4.50 per pound. How many pounds of each type of tea are required?
$5 tea ------------------ $3 tea--------------- mix
x-------------------------(100-x)---------------100 quantities
5x+3*(100-x)= 100*4.5
5x+300-3x=450
2x= 150
x=75 pounds $5 tea
$3 tea = 25 pounds
enjoy your maths